# Magnetocaloric effect for a $Q$-clock type system

### v1

@@ III Mean Field Approximation
The mean-field theory assumes that the fluctuations around the average value of the order parameter, in this case, the magnetization per site $\vec{m}=\vec{M}/N$, are so small that they can be neglected. The first term of the Hamiltonian of Eq. (1) that corresponds to the interaction term between the spin of the lattice in different sites is modified by performing the following approximations.
@@
where now the sum runs for each site in the lattice, and $z$ are the effective nearest neighbors of the model. Consequently, we can define a Hamiltonian per site given by the structure
@@
Amplifying the above qualitatively, we propose an expression for the number of near neighbors effective that adjusts according to the weighting of the effect of non-interacting edges in the system when the lattice has a generic resolution $L\times L$. The effective neighbors for a central spin for the mean-field are determined by the following expression (independent of $Q$) for the square lattice with up, down, left, and right near neighbors [43]
@@
IV Thermodynamic Quantities
Once the partition function of the system has been obtained by either of the two approaches discussed above, it is possible to compute all the thermodynamical observables in a general way through the expressions ( with $k_{B}=1$ )
@@ IV Thermodynamic Quantities
$U=T^{2}\frac{\partial\ln Z}{\partial T},$ (10)
@@
$C=\frac{\partial U}{\partial T},$ (11)
@@
where $F$ is the Helmholtz free energy, $U$ is the internal energy and $\mathcal{C}$ is the specific heat at constant magnetic field. In addition, with the differential expression of Helmholtz free energy given by $dF=-SdT-MdB$, we can obtain the entropy and the magnetization of the system given by
$S=-\frac{\partial F}{\partial T};\quad M=-\frac{\partial F}{\partial B}.$ (12)
Figure 2: Specific heat as a function of temperature for even $Q$ values between $Q=2$ and $Q=8$ for (a) $B=0.05$ and (b) $B=0.1$ for a $3\times 3$ lattice. We observe the peak displacements to the right at a higher magnetic field.
We will begin by reviewing the thermodynamic results of the model, focusing first on the analysis of the specific heat for the exact results of a $3\times 3$ lattice. This thermodynamic quantity for a field $B=0.05$ and field $B=1$ is shown in Fig. 2(a) and Fig. 2(b), respectively for $Q=2,4,6,8$. Here, we observe two peaks for $q\geq 6$, indicative of a double-phase transition. It is, therefore, the BKT phase that appears in these cases. It is also noticeable that the critical temperature of each transition increases with the increasing external field, as we can observe from Fig. 2(b) compared with the critical temperatures obtained for (a) panel of the same figure. This is because the external field favors ordered phases (FM and BKT) over disordered ones. A different behavior is obtained by fixing the field and increasing $Q$, where we observe that the high-temperature peak remains at the same temperature, whereas the low-temperature peak tends toward lower temperatures (eventually to a zero temperature) as $Q$ increases, tending to infinity. Therefore, for $Q\rightarrow\infty$, we expect the temperature of the first transition (from FM to BKT phase) to go to zero. The case of internal energy is shown in Fig. 3(a), where we observe that we obtain lower values for the case with a magnetic field than for the case without an external magnetic field. This is due to the Zeeman term present in the Hamiltonian given by Eq. (1), which decreases the ground state energy compared to the case in the absence of a magnetic field. Continuing with the thermodynamic description of our system of study, Fig. 3(b) shows the magnetization of the system for $Q=2,4,6,$ and $8$ for high fields $B=1$ (blue curves) and $B=0.05$ (purple curves). First, we notice that all systems start saturated (in a ferromagnetic state) and then decrease their magnetization as the temperature increases. The case with a higher external magnetic field shows higher magnetization than the case with a lower magnetic field as the temperature increases for all $Q$ displayed. It is observed that the change in magnetization as a function of temperature happens much faster for larger and larger q values. This is a direct consequence of the degrees of freedom of the system. For example, we can think of $Q=2$, where two possible orientations are up and down. In terms of energy, generating a flip of such a configuration is much more difficult. Therefore, the magnetization at the same temperature will be higher than in the other cases for higher values of $Q$.
Figure 3: Internal energy (a), magnetization (b), and entropy (c) as a function of temperature for different values of $Q$ pairs between $Q=2$ and $Q=8$ for fields $B=0.05$ (purple) and $B=0.1$ (blue) for a $3\times 3$ lattice.
Next, we analyze the entropy in Fig. 3(c). We observe that for a given $Q$, the entropy is lower for the case with field $B=1$ than for the case of $B=0.05$. This is consistent with the interpretation that the magnetic field tends to order the spin system. Also, it is noticeable that for more spin degrees of freedom (as $Q$ increases), the entropy is higher as a function of temperature consistent with the greater number of accessible states the system acquires. These graphs show that the entropy fulfills that $\lim_{T\rightarrow 0}S(T,B\neq 0)=0$ because, with even a very low field value, the system at low temperature tends to have only one preferred state due to the presence of the external field applied to the system. It is important to remember that in the case of a null field, the entropy at low temperature has the value of $\lim_{T\rightarrow 0}S(T,0)=\ln Q$.
Figure 4: In panel (a), we show the spin magnetization of a $3\times 3$ lattice with an external magnetic field equal to unity for different values of $Q=2,4,6,8$ calculated exactly (orange color) and approximated via mean field (blue color). In panel (b), we exhibit the spin magnetization of a $3\times 3$ lattice with an external magnetic field close to zero ($B=0.05$ in $\mathcal{J}$ units) for different values of $Q=2,4,6,8$ calculated exactly (purple color) and approximated via mean field (green color).
In Fig. 4(a), we note the saturation in the magnetization for low temperatures, finding for all the cases studied the unity for the limit of $T\rightarrow 0$. Additionally, we observe that the temperature increase increases the difference between the mean field and the exact calculation for the Ising case $(Q=2)$. In contrast, for $Q\geq 4$, the difference between both methods decreases as the temperature increases. In Fig. 4(b), we find the spin magnetization of a $3\times 3$ lattice with an external magnetic field close to zero ($B=0.05$ in $\mathcal{J}$ units) for different values of $Q=2,4,6,8$ calculated in an exact way (purple color) and approximated via mean field (green color). We note a difference in the shape of the curves as the temperature increases, although both methods start to become saturated at low temperatures. The shape of the exact calculated curve shows that the magnetization behavior for low non-zero external fields varies with increasing field and not so for the medium field case.
V Caloric Phenomena
To analyze this effect, we start by treating the total entropy of a material as the sum of three entropies: electronic $S_{e}$, magnetic $S_{m}$, and lattice $S_{l}$.
$S(T,B)=S_{e}(T,B)+S_{m}(T,B)+S_{l}(T,B)$ (13)
Magnetic entropy depends strongly on the magnetic field, whereas it is very common to find materials with electronic and lattice entropy with negligible dependence on the magnetic field. However, this is not general, and we highlight that in a low temperatures regime ($\approx$<≈ < 10 K), some reports indicate that the electronic entropy can present nonlinear dependence on the magnetic field. In this work, let us assume that we are dealing with a more usual system, where we can establish null dependence of the magnetic field in the lattice and electronic entropies so that Eq. (13) is of the form:
$S(T,B)\approx S_{e}(T)+S_{m}(T,B)+S_{l}(T)$ (14)
The quantification of the caloric phenomenon is associated with a thermodynamic refrigerator cycle, where we can take two paths to quantify the effect: i) an adiabatic trajectory and ii) an isothermal trajectory. The temperature variation that suffers the systems along the isentropic process is called $\Delta T_{ad}$ and is given by the expression
[…]
where we use that $C_{B}=\left(\frac{\partial U}{\partial T}\right)_{B}=T\left(\frac{\partial S}{\partial T}\right)_{B}$ corresponds to specific heat at constant $B$ parameter.
In the case of quantifying the effect employing an isothermal trajectory, we obtain an entropy variation at constant temperature ($\Delta S$) given by
[…]
If we look at Eq. (15) and Eq. (16), we can find a relationship between these quantities concerning each other. It is found that $-\Delta S\propto\Delta T_{ad}$. Consequently, it is essential to note that when you have a case in which $-\Delta S>0$, let’s call this kind of response of direct type. The system will heat up, while when the response is of the $-\Delta S<0$ type, we call this inverse type response, and consequently, the system will cool down. Therefore, we expect for direct response a $\Delta T_{ad}>0$, and for the case of an inverse response, we expect a $\Delta T_{ad}<0$ for the final result of the caloric phenomena.
For the case of the entropy variation at constant temperature ($T$), the expression can be given by the difference of the entropy at the initial and final point of the process in the following form:
$-\Delta S=-\Delta S_{m}=S_{m}(T,B_{0})-S_{m}(T,B),$ (17)
where the contributions of $S_{l}$ and $S_{e}$ have been canceled due to its dependence only on temperature, and consequently, in an isothermal process, it will have no associated variations. The last equation generates a graph for a given $B_{0}$ and final $B$ as a function of $T_{0}$ that quantifies the caloric phenomena.
VI Results
Figure 5: (a) Entropy as a function of temperature for the case $Q=2$ for the case of $B=0.05$ (purple curve), $B=0.5$ (green curve) and $B=1$ (blue curve). (b) Entropy difference $-\Delta S=S(T,B_{0}=0.05)-S(T,B)$ as a function of temperature for $B=0.5$ (green curve), $B=1$ (blue curve) and $B=0.05$ (purple curve).
To start the discussion of the caloric phenomena, we must fix the zero value for temperature and the external magnetic field for our calculations. The zero temperature for our thermodynamic quantities will be $T=0.01$ while the smallest field we will work with will be $B=0.05$. These two quantities combined produce zero entropy, as theoretically expected. In addition, the selection of $B=0.05$ as the lowest field is supported to avoid the problem generated when we have $B\rightarrow 0$ in the mean-field self-consistency equation. Regarding the maxima used in our calculations, the maximum temperature corresponds to $T=3.01$ and $B=1$ in units of $\mathcal{J}$, both quantities.
VI.1 Small Lattices
Let us start the analysis with a simple example of the exact case ($3\times 3$ lattice) for $Q=2$ to understand the caloric response of the system. For this purpose, we plotted the entropy for three different field values of the system in Fig. 5(a), where $B=0.05$ is the purple curve, $B=0.5$ is the green curve, and B=1 is the blue curve. The entropy difference that quantifies the effect is given by $-\Delta S=S(T,0.05)-S(T,B)$, where we have selected the initial point with an external field strength $B=0.05$. These entropy differences can be plotted for different field values on a curve as a function of the temperature to which the process is taken, and they are presented in Fig. 5(b). From this last figure, we note that the response of the effect is positive, which indicates a direct type of behavior for both cases presented. It is also observed that below 0.9 temperature (approximately), the thermal response of the effect for $B=0.5$ and $B=1$ are equal. After this temperature value, it is observed that the case with the higher field is notably different from the one with a lower field. The explanation of this phenomenon is relatively simple. The system prefers to be in a ferromagnetic configuration in an external magnetic field, i.e., in a single configuration state. To remove it from this state requires a considerable increase in temperature in terms of energy. This is reflected in the entropy for cases $B=0.5$ (green curve) and $B=1$ (blue curve) in Fig. 5(a) where both entropies are equal up to $T\sim 0.9$ and then differentiate with $B=0.5$ being higher than $B=1.0$.
Figure 6: Caloric response with an initial $B$ of $0.05$ for the case of a final field of (a) $B=0.1$ and (b) $B=1$ for exact calculations for a $3\times 3$ lattice and using the mean-field approximation.
To compare the effects for the exact calculations and mean field, we have analyzed $-\Delta S$ for the case of a final field of $B=0.1$ and $B=1$ and the initial fixed value of field $B=0.05$. From Fig. 6(a), we note a large discrepancy in the effect’s location as a temperature function for the case with a final $B$ of $0.1$. As we observe in Fig. 6(b), this discrepancy decreases as we increase the final field. It is remarkable from these two comparative plots that if we only focus on estimating the maximum effect at the value associated only with $-\Delta S$ and not the location of the temperature, both calculations do not differ too much. We will return to this point later in the discussion when we deal with other $Q$ cases.
@@ VI.1 Small Lattices
We will now analyze the MCE for different values of $Q$ of the $3\times 3$ lattice, both in the exact case and using the mean field. This is presented in Fig. 7, where the cases for a final field of $B=0.1$ (panel (a)) and the case of $B=1$ (panel (b)) with an initial field of $B=0.05$ are shown. We first notice that for the value of $Q=4$, what was observed for $Q=2$, associated with the location of the maximum temperature effect for both methods, is also true. As the magnetic field increases, the maximum temperature locations become closer. However, a considerable error is obtained when extrapolating mean-field solutions for small lattices, and it has to do with the location of the maximum (or maxima) at $-\Delta S$ when we have the cases $Q=6$ and $Q=8$ as we can appreciate in Fig. 7(b). From this last figure, it can be seen that both $Q=6$ and $Q=8$ present two maxima in $-\Delta S$, vestiges of the double transition present for those values of $Q$ in the model, both for the mean field and the exact calculations. We note that the exact calculations indicate that the peak at the lowest temperature is where the greatest effect occurs, while the mean field indicates the opposite; the second peak indicates the maximum value of the effect. Our previous work [41, 42] shows this phenomenon in the specific heat up to lattice of approximately $7\times 7$. After that value, the second peak observed in the specific heat becomes larger, following the trend shown and observed in the mean-field approximation. In addition, from Fig. 7(a), we observe that in the exact case for the same parameter and lattice configuration, the maximum of the effect is slightly higher for $Q=6$ and $8$ than for $Q=4$. Meanwhile, the mean-field indicates that the ordering of the effect from highest to lowest is $Q=4$, $Q=6$, $Q=8$, and finally, $Q=2$. In contrast, from Fig. 7(b), we can see that the largest effect is obtained for $Q=4$ independent of the calculation method. This is remarkable since previous studies found that when applying thermodynamic cycles using the $Q$-clock model as the working substance, the maximum efficiency and useful work are obtained for the parameter $Q=4$ [43]. These two cases, therefore, indicate that there is a field value at which this inversion observed in the effect maxima occurs. For this purpose, we plot in Fig. 8 the value of the maximum $-\Delta S_{max}$ as a function of the field independent of the location in temperature at which it occurs both for the mean field and for the exact calculation. From this plot, we can notice that in the exact case, $Q=6$ and $Q=8$, responses higher than $Q=4$ and $Q=2$ are observed up to a field of approximately $\sim 0.26$. After that value, the effect is ordered in the same way as the mean-field results, i.e., in the form:
@@
From Fig. 8, we can also confirm what happens when estimating the maximum effect between the exact and mean-field calculations. Although the mean-field and exact differ quite a bit in the temperature location of the maximum point, if we focus only on the numerical value of effect maximization at $-\Delta S$, the mean-field does not show significant variations compared to the exact calculation even for a small lattice, as shown in this study. The dotted curves in Fig. 8 representing mean-field calculations differ very little from the solid lines of the exact case, with the largest difference observed for $Q=4$ and the smallest differences for $Q=6$ and $Q=8$. From this tendency, we can conclude that for values larger than $Q=4$, the numerical estimate of $-\Delta S$ will be closer between the two methods presented. We can say, then, that the mean field is useful in this context even for small lattices in the numerical estimation of the maximum effect but fails logically in the location of the temperature where it occurs because the nature of the approximation used is regularly employed for lattices of larger sizes.
@@
An interesting point to discuss is the location in temperature where the maximum value of the effect is given compared to the critical temperature $(T_{c})$ of the different phases present in the model. This is shown in Fig. 9, which represents a plot of $T$ vs. $B$ for (a) $Q=2$ and $Q=4$ and for (b) $Q=6$ and $Q=8$. To understand these plots, we will start with the cases $Q=2$ and $Q=4$ presented in Fig. 9(a). We must think that the variation of the effect is given by $-\Delta S=S(T,B_{0}=0.05)-S(T,B)$ and hence is described by a spin system starting with an entropy of the form $S(T,B_{0}=0.05)$. This quantity has associated a specific heat $C(T,B_{0}={0.05})$ that defines a critical temperature represented by horizontal lines in Fig 9(a). On the other hand, the continuous lines represent the final entropy state $S(T,B)$, which has an associated specific heat $C(T,B)$, which will deliver as the field changes different critical temperature values, indicating the respective phase transition for these cases. The graphs with geometrical figures (triangular for $Q=2$ and square for $Q=4$) are the points where, for a given field, $-\Delta S$ is the maximum. Precisely, the intersection between the curve with geometric figures and the dotted horizontal one generates, for $Q=2$ and $Q=4$, two possible regions where the system maximizes its effect. Under a field value given by the vertical black line, the system maximizes its effect in a region where the system starts in a ferromagnetic phase $(FP_{B_{0}})$ and ends in a ferromagnetic phase ($FP_{B}$) while past that critical field value found the system maximizes the caloric response starting from a paramagnetic phase ($PP_{B_{0}}$) and ending in a ferromagnetic phase ($FP_{B}$).
The same analysis can be made if we recall that for the exact case of $Q=6$ and $Q=8$, the peak of the effect was concentrated in the first peak of Fig. 7 and, therefore, more oriented to transitions relating to the BKT and FP phases. This is shown in Fig. 9(b), where the maxima are concentrated when the system remains in a ferromagnetic phase or involves transitions linking the BKT and FP phases. We note that the maxima of the effect for $Q=8$ are much closer to the $T_{c}$ of the first phase transition related to an FP-BKT type transition of the final state of the spin system than what occurs for the case of $Q=6$.
VI.2 Large Lattices
Including mean-field calculations allows for studying the caloric response for larger lattice sizes and analyzing the convergence of results. Fig. 10(a) presents the spin caloric response as a function of temperature for $Q=4$, and in Fig.(b) for the case of $Q=8$ both for lattice sizes of $3\times 3$, $8\times 8$, $16\times 16$ and $32\times 32$ with an initial field of $B=0.05$ and a final field of $B=1$. As discussed, the $Q=4$ case only presents one transition of type FP to PP, and we therefore see only a peak in the entropy difference, while $Q=8$ presents two phase transitions where we see that $-\Delta S$ is larger when we go from BKT to PP over that from FP to BKT. From both graphs, we notice that there is a convergence of the caloric response already for a $32\times 32$ lattice because it presents the same results as a $16\times 16$ lattice both for the case of $Q=4$ and $Q=8$. It can also be observed by comparing the two graphs that it is still $Q=4$ that presents the highest $-\Delta S/N$ evaluated on the system.
Figure 10: Mean field calculations for (a) $Q=4$ and (b) $Q=8$ for $3\times 3$, $8\times 8$, $16\times 16$ and $32\times 32$ lattice.
Figure 11: Magneto-caloric per spin effect for $16\times 16$ and $32\times 32$ lattices using mean-field calculations (orange curves) and using the Monte Carlo method (black-grey curves) for the case (a) $Q=4$ and (b) $Q=8$.
For the completeness of our calculations, we performed a large lattice scan where the thermodynamic properties of the magnetic substance are obtained through the Monte Carlo method and state selection by the Metropolis algorithm, thus obtaining the internal energy and magnetization for fixed external fields. Fixing the external magnetic field of the simulation at $0.05$ and $1.00$ (in units of J) and calculating the observables with a thermalization of $50000$ flips and a set of $100000$ iterations for the calculation of thermal averages, the curve of the magnetocaloric effect is obtained for $L=16$ and $L=32$. In Fig. 11, we observe the magneto-caloric per spin effect for $16\times 16$ and $32\times 32$ lattices using mean-field calculations (orange curves) and using the Monte Carlo method (black-grey curves) for the case (a) $Q=4$ and (b) $Q=8$. We can observe from there that the Monte Carlo calculation concerning the mean-field calculation coincides with each other in the maximum value of the effect but not in the temperature at which the maximum occurs. It can also be seen that for the case of the Monte Carlo analysis, the $16\times 16$ and $32\times 32$ effects still show differences with each other (small), while, as we mentioned earlier in the discussions, the mean-field case already converges for a $32\times 32$ lattice. However, it is observed that for Monte Carlo simulations up to the location of the first maximum (or the only one for $Q<5$) corresponding to an FP to PP transition, $-\Delta S$ presents almost the same values for $16\times 16$ and $32\times 32$. However, there is a small difference between the values obtained for these networks near the single transition for Q=4 and from the second transition for $Q=8$.
Finally, returning to the discussion of the phases and their role in quantifying the effect, Fig. 11(a)-(b) shows that for large networks, the peak of the effect is much closer to the BKT-PP type transition, and this is corroborated by both mean-field and Monte Carlo simulations. However, we must reiterate that the opposite occurs in small lattices, where exact calculations have been performed since we observe that the maximum $-\Delta S$ is close to the FP-BKT type transitions.
VII Conclusions
In this work, we have studied the magneto-caloric effect for a spin system with $Q$ orientations known as the $Q$-clock model. We reviewed the thermodynamic quantities describing the model and calculated for $Q$ pairs between $2$ and $8$ using three methods: exact, mean-field, and Monte Carlo simulations. For small lattices ($3\times 3$) by exact calculations, we have found that there is a region of field where the effects of $Q=6$ and $Q=8$ are higher than those presented for the cases of $Q=4$ and $Q=2$, and then in higher regimes of the external field, they are ordered in descending form as $-\Delta S_{max}^{Q=4}>-\Delta S_{max}^{Q=6}>-\Delta S_{max}^{Q=8}>-\Delta S_{max}^{Q=2}$. Forcing the mean-field approximation to explore these lattice sizes, we note that mean-field allows us to quantify the maximum (in intensity) fairly close to the exact calculation; however, for reasons of applying it to small lattices, it fails considerably in the temperature location where it occurs.
Related to the magnetic phases involved in maximizing the caloric phenomenon, we can report that for small lattices for $Q\geq 5$, the FP-BKT transition plays a fundamental role. This changes abruptly when we explore large lattices, where Monte Carlo simulations are employed, which indicate that the transition where the effect is maximized is closer to a BKT-PP type. In addition, there are two considerable effects for large-size lattices: First, there is a convergence of results, evidenced by Monte Carlo and mean-field simulations for lattices as large as $32\times 32$ in the results, and the superior effect always occurs for $Q=4$.
VIII Acknowledgments
M. A., F. J. P., E. V., and PV acknowledge the financial support of CEDENNA (Chile) under contract ANID AFB220001 and Fondecyt (Chile) under contract 1230055. F. J. P. and P.V. acknowledge support from ANID Fondecyt grant under contract 1240582. F.J.P. acknowledges financial support from “Millennium Nucleus in NanoBioPhysics” project NNBP NCN2021 _021 and UTFSM DGIIE.

### v2

@@ Appendix A Exact and Mean Field Approximation
The second corresponds to the mean-field theory that assumes that the fluctuations around the average value of the order parameter, in this case, the magnetization per site $\vec{m}=\vec{M}/N$, are so small that they can be neglected. The first term of the Hamiltonian of Eq. (1) that corresponds to the interaction term between the spin of the lattice in different sites is modified by performing the following approximations.
@@
where now the sum runs for each site in the lattice, and $z$ are the effective nearest neighbors of the model. Consequently, we can define a Hamiltonian per site given by the structure:
@@
Amplifying the above qualitatively, we propose an expression for the number of near neighbors effective that adjusts according to the weighting of the effect of non-interacting edges in the system when the lattice has a generic resolution $L\times L$. The effective neighbors for a central spin for the mean field are determined by the following expression (independent of $Q$) for the square lattice with up, down, left, and right near neighbors [43]
@@
Once the partition function of the system has been obtained by either of the two approaches discussed above, it is possible to compute all the thermodynamical observables in a general way through the expressions ( with $k_{B}=1$ ):
@@
$U=T^{2}\left(\frac{\partial\ln Z}{\partial T}\right)_{B},$ (21)
@@
$C=\left(\frac{\partial U}{\partial T}\right)_{B},$ (22)
@@
where $F$ is the Helmholtz free energy, $U$ is the internal energy, and $C$ is the specific heat at the constant magnetic field. In addition, with the differential expression of Helmholtz free energy given by $dF=-SdT-MdB$, we can obtain the entropy and the magnetization of the system given by
$S=-\left(\frac{\partial F}{\partial T}\right)_{B};\quad M=-\left(\frac{\partial F}{\partial B}\right)_{T}.$ (23)
@@
A.1 Small Lattices
We will now analyze the MCE for different values of $Q$ of the $3\times 3$ lattice, both in the exact case and using the mean field. This is presented in Fig. 7, where the cases for a final field of $B=0.1$ (panel (a)) and the case of $B=1$ (panel (b)) with an initial field of $B=0.05$ are shown. We first notice that for the value of $Q=4$, what was observed for $Q=2$, associated with the location of the maximum temperature effect for both methods, is also true. As the magnetic field increases, the maximum temperature locations become closer. However, a considerable error is obtained when extrapolating mean-field solutions for small lattices, and it has to do with the location of the maximum (or maxima) at $-\Delta S$ when we have the cases $Q=6$ and $Q=8$ as we can appreciate in Fig. 7(b). From this last figure, it can be seen that both $Q=6$ and $Q=8$ present two maxima in $-\Delta S$, vestiges of the double transition present for those values of $Q$ in the model, both for the mean field and the exact calculations. We note that the exact calculations indicate that the peak at the lowest temperature is where the most remarkable effect occurs, while the mean field suggests the opposite; the second peak indicates the maximum value of the effect. In addition, from Fig. 7(a), we observe that in the exact case for the same parameter and lattice configuration, the maximum of the effect is slightly higher for $Q=6$ and $8$ than for $Q=4$. Meanwhile, the mean field indicates that the ordering of the effect from highest to lowest is $Q=4$, $Q=6$, $Q=8$, and finally, $Q=2$. In contrast, from Fig. 7(b), we can see that the most prominent effect is obtained for $Q=4$ independent of the calculation method. This is remarkable since previous studies found that when applying thermodynamic cycles using the $Q$-clock model as the working system, the maximum efficiency and useful work are obtained for the parameter $Q=4$ [43]. These two cases, therefore, indicate that there is a field value at which this inversion observed in the effect maxima occurs. For this purpose, we plot in Fig. 8 the value of the maximum $-\Delta S_{max}$ as a function of the field independent of the location in temperature at which it occurs both for the mean field and for the exact calculation. From this plot, we can notice that in the exact case, $Q=6$ and $Q=8$, responses higher than $Q=4$ and $Q=2$ are observed up to a field of approximately $\sim 0.26$. After that value, the effect is ordered in the same way as the mean-field results, i.e., in the form:
@@ A.1 Small Lattices
From Fig. 8, we can also confirm what happens when estimating the maximum effect between the exact and mean-field calculations. Although the mean-field and exact differ quite a bit in the temperature location of the maximum point, if we focus only on the numerical value of effect maximization at $-\Delta S$, the mean-field does not show significant variations compared to the exact calculation even for a small lattice, as shown in this study. The dotted curves in Fig. 8 representing mean-field calculations differ very little from the solid lines of the exact case, with the most considerable difference observed for $Q=4$ and the minor discrepancies for $Q=6$ and $Q=8$. From this tendency, we can conclude that for values larger than $Q=4$, the numerical estimate of $-\Delta S$ will be closer between the two methods presented. We can say that the mean field is helpful in this context, even for small lattices, in the numerical estimation of the maximum effect. However, it fails logically in the location of the temperature where it occurs because the nature of the approximation used is regularly employed for lattices of larger sizes.
@@
An interesting point to discuss is the location in temperature where the maximum value of the effect is given compared to the critical temperature $(T_{c})$ of the different phases present in the model. This is shown in Fig. 9, which represents a plot of $T$ vs. $B$ for (a) $Q=2$ and $Q=4$ and for (b) $Q=6$ and $Q=8$. To understand these plots, we will start with the cases $Q=2$ and $Q=4$ presented in Fig. 9(a). We must think that the variation of the effect is given by $-\Delta S=S(T,B_{0}=0.05)-S(T,B)$ and hence is described by a spin system starting with an entropy of the form $S(T,B_{0}=0.05)$. This quantity has associated a specific heat $C(T,B_{0}={0.05})$ that defines a critical temperature represented by horizontal lines in Fig 9(a). On the other hand, the continuous lines represent the final entropy state $S(T,B)$, which has an associated specific heat $C(T,B)$, which will deliver as the field changes different critical temperature values, indicating the respective phase transition for these cases. The graphs with geometrical figures (triangular for $Q=2$ and square for $Q=4$) are the points where, for a given field, $-\Delta S$ is the maximum. Precisely, the intersection between the curve with geometric figures and the dotted horizontal one generates, for $Q=2$ and $Q=4$, two possible regions where the system maximizes its effect. Under a field value given by the vertical black line, the system maximizes its effect in a region where the system starts in a ferromagnetic phase $(FP_{B_{0}})$ and ends in a ferromagnetic phase ($FP_{B}$). In contrast, past that critical field value found the system maximizes the caloric response starting from a paramagnetic phase ($PP_{B_{0}}$) and ending in a ferromagnetic phase ($FP_{B}$). The same analysis can be made if we recall that for the exact case of $Q=6$ and $Q=8$, the peak of the effect was concentrated in the first peak of Fig. 7 and, therefore, more oriented to transitions relating to the BKT and FP phases. This is shown in Fig. 9(b), where the maxima are concentrated when the system remains in a ferromagnetic phase or involves transitions linking the BKT and FP phases. We note that the maxima of the effect for $Q=8$ are much closer to the $T_{c}$ of the first phase transition related to an FP-BKT type transition of the final state of the spin system than what occurs for the case of $Q=6$.